Minute of KAGRA commissioning meeting
- Date: 2018/11/20 11AM~12PM
- Attendance: Miyakawa, Aso, Enomoto, Izumi, Lucia
- Zoom: Kozakai, Yamamoto, Koseyama, Oshino, Suzuki, Yokozawa, Ushiba
- CAL injection on 11/26-29 won't affect to commissioning task.
- We need 2weeks more.
- Minimum success is stabilization of IR fluctuation to cavity resonant frequency.
- X arm lock
- 2hours lock was lost not by instability of VIS, by something but we are not sure.
- IR flashes could be seen with CARM lock, but it is not sure it was TEM00. We could know FSR of cavity by the measurement.
- Yarm commissioning
- Nakano-kun, Eleonora lead the Y arm commissioning task, and Sugimoto-kun can join to Y arm team. Enzo and Type B team can work for commissioning work specially for BS and SR.
- Laser, PSL, IOO report
- Characterization for PMC, IMC, FSS started. The results will be summerized to somewhere.
- ISS noise hunting started. No difference was seen between KIKUSUI and KEPCO DC power supply.
- Next meeting: 11/27.